OpenStudy (anonymous):

Find x if y = –1. 6 – 3xy = 12 A.1/2 B.2 2 C.-1/2 D. –2

OpenStudy (uri):

6 – 3xy = 12 6-3x(-1)=12 6+3x=12 12-6=3x 6=3x x=6/3 x=2